Circle the correct answer:

Let:     A = number of units of product A produced on Saturday,

            B = number of units of product A produced on Sunday,

            C = number of units of product B produced on Saturday, and

            D = number of units of product B produced on Sunday.

a. Which constraint(s) say that a total of 200 units of product A will be produced on Saturday and Sunday (together).

  1. A= 200   and   B= 200
  2. A+ B= 200
  3. A+ B + C + D= 200
  4. A+ C= 200
  5. 200A= B
  6. None of the above.

b. Which of the following constraints says that half as much of product A will be produced on Saturday as on Sunday?

  1. A <= 2B
  2. B <= 2A
  3. A = 2B
  4. B = 2A
  5. None of the above.
Answer #1

a) A+B = 200

The below table summarizes the given details

Product A Product B Saturday Sunday 0

For this question we need a constraint that states that a total of 200 units of product A are produced on Saturday and Sunday (together). Since we are only asked about Product A, option C and D are not correct. We can see from above table that total units of Product A produced are the sum of A and B. We need that this total should be 200. Hence the constraint is A+B=200. Hence the correct option is option B. Option A is wrong as this states that 200 units each a=of Product A are produced on Saturday and Sunday. Option E is wrong as it states that units of product A produced on Sunday is 200 times that produced on Saturday.

b) B = 2A

It is required that units of product A produced on Saturday is half of units produced on Sunday. From the above table we can see that units of product A produced on Saturday is A and units of product A produced on Sunday is B. We need that A should be equal to half of B. Hence we the below

A = 0.5B = (1/2)B

2A = B

B = 2A

Thus option d is the correct answer.